3. Ow I Stuffed Animals wants to measure the efficiency of its raw materials inventory management. Based on a beginning inventory of $3,000 and an ending inventory of $4,000 and using $48,000 worth of raw materials, what is the days sales in raw materials inventory? a. 22.81 days b. 26.61 days c. 30.42 days d. 52.23 days
Answer #1
Average raw materials inventory = (3000+4000)/2=$3500
Days' sales in raw materials inventory = 365/Materials used*Average inventory
Days' sales in raw materials inventory = 365/48000*3500= 26.61 days
Option B is correct